OpenStudy (anonymous):

If f (x)=3x-9 and g (x)=x^2, what is (g°f)(5)? PLEASE HELP

jigglypuff314 (jigglypuff314):

(g*f)(x) = g(x) * f(x) = x^2 * (3x - 9) = ?

jigglypuff314 (jigglypuff314):

then plug in 5 for x